OpenStudy (anonymous):

Use addition or substitution to solve this problem. To paint his house, Samuel bought 2 cans of paint and one roller for $30. When friends came over to help, he went back to the store and bought three more cans of paint and two more rollers for $50. How much did he pay for a can of paint?

OpenStudy (anonymous):

Soo we have : 2x+y=30 3x+2y=50 From the first equation you have : y=30-2x you substitute y in the second equation and you get : 3x+2.(30-2x) =50 3x+60-4x=50 -x=-10 or x=10 Soo, he paid 10$ for a can of paint